# Normalization in Attention Dynamics

### v1

@@ Abstract
We study the effect of normalization schemes on token representations in deep transformers. Modeling their evolution as interacting particles on the sphere, we show that normalization acts as a form of speed regulation. This perspective enables a unified analysis of several schemes—including Post-LN, Pre-LN, Mix-LN, Peri-LN, nGPT, and LN-Scaling—revealing how they influence clustering dynamics and representation collapse. Our framework clarifies how different schemes shape token representations across layers and provides a principled basis for comparing them, identifying Peri-LN as a particularly effective choice.
@@ 1  Introduction
Transformer architectures have revolutionized natural language processing and beyond, demonstrating unprecedented performance across diverse tasks—from machine translation and text generation to reasoning and protein folding. The remarkable capabilities of transformers, including their emerging reasoning abilities, are enabled by the attention mechanism introduced in Bahdanau et al., (2015); Vaswani et al., (2017).
@@
A recent line of theoretical work, initiated in Geshkovski et al., (2023), studies information processing across deep transformer layers by reframing them as interacting particle systems, building on the original setup of Sander et al., (2022). Following this initial work, layer normalization (LayerNorm) emerged as a critical component significantly influencing the long-term dynamics of these systems. Geshkovski et al., (2025) proposed a model in which particles are constrained to evolve on a sphere, corresponding to the so-called Post-layer norm (Post-LN) scheme. This model has since become the standard paradigm for transformer analysis in subsequent research (Karagodin et al.,, 2024; Geshkovski et al., 2024a, ; Geshkovski et al., 2024b, ; Bruno et al., 2025a, ; Bruno et al., 2025b, ; Criscitiello et al.,, 2024).
@@
Several alternatives to Post-LayerNorm (Post-LN) have emerged in recent years to improve training performance, each subtly altering transformers’ long-term clustering behavior. Most notably, Pre-LayerNorm (Pre-LN) has become the default choice for leading large language models including GPT (Radford et al.,, 2019) and LLaMA (Touvron, H. et al,, 2023). This approach was originally introduced in ResNet-v2 He et al., (2016) before being adapted for Transformer architectures. It enables more stable training of deeper networks while reducing sensitivity to hyperparameters such as learning rates (Xiong et al.,, 2020).
@@
Understanding normalization schemes is essential for advancing the design and performance of transformer architectures. In particular, Sun et al., (2025) and Gromov et al., (2025) identify a phenomenon known as the curse of depth, in which deep layers of large language models (LLMs) degenerate into near-identity transformations. This effect is so pronounced that it enables pruning of deeper layers with minimal impact on performance (Muralidharan et al.,, 2024; Siddiqui et al.,, 2024). On the other hand, the well-known issue of representation collapse presents a significant challenge to increasing the depth of LLMs.
@@
To mitigate this issue, Li et al., (2025) propose a hybrid normalization scheme that applies Post-LN normalization in the early layers and reserves Pre-LN normalization for the deeper layers. This strategy was further refined in the development of Peri-LN (Kim, B., Johnson, M. et al.,, 2025), which has been reported to be used in the Gemma-3 model (Gemma Team et al.,, 2025). Alternatively, Sun et al., (2025) suggest a simpler fix: rescaling Pre-LN by the square root of the depth (a scheme called LN-Scaling). Additionally, Loshchilov et al., (2025) show that with careful architectural design, as in nGPT, normalizing tokens to lie on the unit sphere can further streamline the normalization process.
@@
To answer this question, we revisit both classical and novel LayerNorm schemes through the lens of the simplified interacting particle dynamics introduced in Geshkovski et al., (2025) to bring a theoretical understanding of these various design choices. Since the final decoding layer of a transformer is typically preceded by a normalization step, we focus on the direction of token representations. Regardless of the specific normalization used, these directions naturally form an interacting particle system on the sphere. This shared geometric setting enables a direct, side-by-side comparison of various normalization schemes, all of which we reinterpret as forms of speed regulation. Despite its simplicity, our model captures complex behaviors observed in practice, including curse of depth and representation collapse.
@@
A growing body of work has examined normalization in Transformers, with a primary focus on its empirical and theoretical implications for gradient stability. Notably, Xiong et al., (2020) and Sun et al., (2025) provide experimental evidence that improper placement of normalization layers can lead to exploding or vanishing gradients in deep models. These findings are often supported by variance-based analyses that track the propagation of activations and gradients through the network, such as (Noci et al.,, 2025) and (Kedia et al.,, 2024). Wortsman et al., (2024) further identify normalization-related training instabilities that emerge at scale. Building on this foundation, Li et al., (2025) and Kim, B., Johnson, M. et al., (2025) explore hybrid normalization strategies in large-scale settings, using both theoretical approximations and empirical diagnostics to study gradient flow and the stability of learned representations.
@@
We provide different perspectives on normalization architecture, by casting differently normalized Transformers as variations of a common interacting-particle ODE, where the normalization method determines a speed factor, which can amplify initial velocity and dampen representation collapse in deep layers. Within this unified framework, we extend the framework of Geshkovski et al., (2025) for Post-LN and in particular, establish asymptotic clustering under general conditions on the speed regulation mechanism. To differentiate various normalization schemes we further study the initial and final velocity of tokens corresponding to first and deep layers respectively. In particular, we recover the representation collapse phenomenon that plagues Post-LN. Our theoretical framework identifies Peri-LN as a particularly effective scheme that makes good use of both early and deep layers.
@@ 2.1  Attention
Throughout this paper, we focus on the simplified setting of Geshkovski et al., (2025) where $Q^{t}=K^{t}=V^{t}=I_{d}$ for all $t\geq 0$.
@@
The Root Mean Squared (RMS) norm $\mathsf{Norm}(x)=x/\|x\|$ of a token Zhang and Titov, (2019) is a critical ingredient of all normalization schemes considered here.
@@
In this paper, we study six major schemes: Post-LN (Vaswani et al.,, 2017), Pre-LN (Xiong et al.,, 2020), Mix-LN (Li et al.,, 2025), Peri-LN (Kim, B., Johnson, M. et al.,, 2025), nGPT (Loshchilov et al.,, 2025), and LN-Scaling (Sun et al.,, 2025). Note that Mix-LN is a combination of Post-LN for $t\leq\tau$ and Pre-LN for $t>\tau$ while LayerNorm-Scaling (LN-Scaling) is a deterministic rescaling of Post-LN. The four remaining schemes are presented in Figure 1. Such explicitly layer-normalization rules are not the only strategies employed in practice. Other attempts to improve normalization suggest better initializations $Q^{0},K^{0},V^{0}$ (Kedia et al.,, 2024) and explicit scaling of the updates, similarly to $\alpha_{t}$ that is trainable in nGPT.
@@
LN-Scaling $X^{t+1}=\mathsf{Norm}\bigl(X^{t}+\frac{1}{\sqrt{t+1}}A^{t}(X^{t})\bigr)$ $\dot{X}(t)=\frac{1}{\sqrt{t+1}}\mathbf{P}_{X(t)}A^{t}(X(t))$
@@ 2.3  Speed regulation formulation
In Post-LN, nGPT, and LN-Scaling, tokens are constrained to the the sphere ${\mathcal{S}}^{d-1}$ with LN-Scaling simply adjusting the speed of the particles as a function of $t$ compared to Post-LN. For the other rules where tokens may have varying magnitude, one final projection is typically applied before the final decoding layer in practice. In particular, this means that decoding depends on directions $\theta_{j}(t)=\mathsf{Norm}(x_{j}(t))$, for $j=1,\ldots,n$.
@@ 3  Asymptotic clustering
Since the work of Geshkovski et al., (2023, 2025), theoretical analyses of attention dynamics have primarily focused on establishing asymptotic clustering under the Post-LN scheme, namely $\theta_{j}(t)\to\theta^{*}$ as $t\to\infty$ for all $j=1,\ldots,n$, under a generic initialization; see also Criscitiello et al., (2024); Chen et al., (2025). However, empirical studies have revealed that in practice, tokens often remain trapped in metastable states for extended periods before clustering emerges (Geshkovski et al., 2024a, ; Bruno et al., 2025a, ). Despite this, the clustering phenomenon appears to occur at multiple local scales, and the simplified setting considered in prior work continues to offer valuable insights, as we will demonstrate in the next section. In this section, we extend the analysis and show that asymptotic clustering persists beyond the original Post-LN framework to other normalization schemes.
@@
The main observation of Geshkovski et al., (2025) is that when $KQ^{\top}=QK^{\top}=V$, the Post-LN system is a gradient flow for the energy function
@@
Consider the normalized attention dynamics (NA) with $Q=K=V=I_{d}$. Then for uniformly sampled initializations $\Theta(0)\in({\mathcal{S}}^{d-1})^{\otimes n}$ Post-LN, nGPT, LN-Scaling cluster asymptotically
@@
In fact, this result holds not only for $Q^{t}=K^{t}=V^{t}=I_{d}$ but more generally for $Q^{t}=Q,K^{t}=K$, and $V^{t}=V=Q^{\top}K=K^{\top}Q$ as in Sander et al., (2022). The second condition on the magnitude growth can be traced with Table 2 definition to work with further. For example, we immediately get the following.
@@ 4.1  Prelude: Symmetric initialization
Following Geshkovski et al., (2025); Cowsik et al., (2024), we begin with a so-called orthogonal symmetric initialization where $\langle\theta_{j}(0),\theta_{k}(0)\rangle=0$ for $j\neq k$ and $r_{j}(0)=1$ for all $j$. This configuration approximately matches that of randomly initialized tokens in high dimension. Due to the symmetry, the cosine similarity $\gamma(t)=\langle\theta_{j}(t),\theta_{k}(t)\rangle$ does not depend on $j\neq k$ and the entire token dynamics reduces to the evolution of two scalar quantities: $\gamma(t)$ and $r(t)$. In the Appendix, we derive a simple ODE for $\gamma(t),r(t)$ following Geshkovski et al., (2025, Theorem 6.8). We plot ODE-based evolution of $\gamma(t)$ in Figure 2 with parameters $\beta=5$, $n=256$. Despite its simplicity, this setup already provides striking insight into the effects of different normalization schemes. The importance tracks in how cosine similarity evolution is alike in the theoretical formula plotted in Figure 2 and the experimental setup with random weights modeled in Figure 4.
@@
LN-Scaling $\frac{2}{e^{\beta}+n-1}$ $C\frac{e^{-4\sqrt{t}}}{\sqrt{t}}$
@@
A few remarks are in order. First, the initial velocities are comparable across models, up to the effects of the tuning parameters $\alpha_{0}$ and $r_{0}$. Notably, the temperature parameter $\beta$ exponentially damps the initial velocity, suggesting that initializing $Q$ and $K$ with smaller magnitudes in the early layers may be beneficial. More striking is the effect of speed regulation at terminal velocity: Pre-LN, Mix-LN, and nGPT (with constant $\alpha_{t}$) exhibit a polynomial slowdown, in contrast to other normalization schemes. While LN-Scaling converges more slowly than exponential, it still outpaces the polynomial decay. This implies that Pre-LN, Mix-LN, and nGPT cluster more gradually than their counterparts—indicating a more effective use of intermediate layers and a stronger resistance to representation collapse. Finally, note that the trainable parameter $\alpha_{t}$ in nGPT can have a drastic impact on both initial and terminal velocity. See Figure 2 for a visual representation of cosine similarity and evolution of $\dot{\gamma}$ relative to time and position. See Figure 3 for comparison between different $\alpha_{t}$ in nGPT.
@@ 4.2  Initial velocity
To interpret the significance of Theorem 4.2, recall from Table 2 that the initial velocity of direction $\theta_{j}$ is dampened by a factor proportional to $\|A_{j}(0)\|$ for both Peri-LN and nGPT. Consequently the first–layer angular displacement of Peri-LN and nGPT exceeds that of Post-LN, Pre-LN, Mix-LN, and LN-Scaling, by a factor $\Omega(\min(d/\log n,\sqrt{n/\log n}))$.
@@ 4.3  Terminal velocity
Radial Growth under Pre-LN. Our first goal is to estimate the rate of growth of $r_{j}(t)$, the norm of token $j$’s representation, under Pre-LN normalization. Empirically, the growth of hidden states in transformers has been well-documented. For instance, studies such as (Xiong et al.,, 2020; Kedia et al.,, 2024) observe that in randomly initialized transformers, $r_{j}(t)\sim\sqrt{t}$, reflecting the diffusive nature of a random walk induced by randomly sampled projections $V$.
@@
Again this result corroborates the findings of Section 4.1: Post-LN and LN-Scaling cluster token directions at an exponential rate, while Pre-LN, Peri-LN and Mix-LN slow down to a polynomial ($\sim 1/t^{C}$) decay. Moreover, nGPT has the ability to control rate of clustering through $\alpha_{t}$. This confirms that Pre-LN makes better use of depth, as tokens continue to evolve meaningfully across many layers, rather than collapsing too quickly. In particular, this difference explains why Pre-LN is less prone to representation collapse in very deep models compared to Post-LN.

### v2

@@ Abstract
We study the effect of normalization schemes on token representations in deep transformers. Modeling their evolution as interacting particles on the sphere, we show that normalization acts as a form of speed regulation. This perspective enables a unified analysis of several schemes—including Post-LN, Pre-LN, Mix-LN, Peri-LN, nGPT—revealing how they influence clustering dynamics and representation collapse. Our framework clarifies how different schemes shape token representations across layers and provides a principled basis for comparing them, identifying Peri-LN as a particularly effective choice.
@@ 1  Introduction
Transformer architectures have revolutionized natural language processing and beyond, demonstrating unprecedented performance across diverse tasks—from machine translation and text generation to reasoning and protein folding. The remarkable capabilities of transformers, including their emerging reasoning abilities, are enabled by the attention mechanism introduced in BahChoBen15a; vaswani2017attention.
@@
A recent line of theoretical work, initiated in geshkovski2023emergence, studies information processing across deep transformer layers by reframing them as interacting particle systems, building on the original setup of sander2022sinkformers. Following this initial work, layer normalization (LayerNorm) emerged as a critical component significantly influencing the long-term dynamics of these systems. mathpersp23 proposed a model in which particles are constrained to evolve on a sphere, corresponding to the so-called Post-layer norm (Post-LN) scheme. This model has since become the standard paradigm for transformer analysis in subsequent research (KarPolRig24; GesKouPol24; GesRigRui24; BruPasAga24; bruno2025multiscale; boumal).
@@
Several alternatives to Post-LayerNorm (Post-LN) have emerged in recent years to improve training performance, each subtly altering transformers’ long-term clustering behavior. Most notably, Pre-LayerNorm (Pre-LN) has become the default choice for leading large language models including GPT (radford2019gpt) and LLaMA (touvron2023llama). This approach was originally introduced in ResNet-v2 he2016identity before being adapted for Transformer architectures. It enables more stable training of deeper networks while reducing sensitivity to hyperparameters such as learning rates (xiong2020layer).
@@
Understanding normalization schemes is essential for advancing the design and performance of transformer architectures. In particular, sun2025curse and gromov2024unreasonable identify a phenomenon known as the curse of depth, in which deep layers of large language models (LLMs) degenerate into near-identity transformations. This effect is so pronounced that it enables pruning of deeper layers with minimal impact on performance (muralidharan2024compact; siddiqui2024deeper). On the other hand, the well-known issue of representation collapse presents a significant challenge to increasing the depth of LLMs.
@@
To mitigate this issue, li2024mixln propose a hybrid normalization scheme that applies Post-LN normalization in the early layers and reserves Pre-LN normalization for the deeper layers. This strategy was further refined in the development of Peri-LN (kim2025peri), which has been reported to be used in the Gemma-3 model (gemma2025). Alternatively, NociAnagBig22 suggest a simpler fix: rescaling residuals by the square root of the depth (a method we will call sqrt-scaling). This method with Pre-LN  architecture was then studied in-depth by sun2025curse. Additionally, loshchilov2024ngpt show that with careful architectural design, as in nGPT, normalizing tokens to lie on the unit sphere can further streamline the normalization process.
@@
To answer this question, we revisit both classical and novel LayerNorm schemes through the lens of the simplified interacting particle dynamics introduced in mathpersp23 to bring a theoretical understanding of these various design choices. Since the final decoding layer of a transformer is typically preceded by a normalization step, we focus on the direction of token representations. Regardless of the specific normalization used, these directions naturally form an interacting particle system on the sphere. This shared geometric setting enables a direct, side-by-side comparison of various normalization schemes, all of which we reinterpret as forms of speed regulation. Despite its simplicity, our model captures complex behaviors observed in practice, including curse of depth and representation collapse.
@@
A growing body of work has examined normalization in Transformers, with a primary focus on its empirical and theoretical implications for gradient stability. Notably, xiong2020layer and sun2025curse provide experimental evidence that improper placement of normalization layers can lead to exploding or vanishing gradients in deep models. These findings are often supported by variance-based analyses that track the propagation of activations and gradients through the network, such as (NociAnagBig22) and (kedia2024transformersstableendtoendsignal). wortsman2024small further identify normalization-related training instabilities that emerge at scale. Building on this foundation, li2024mixln and kim2025peri explore hybrid normalization strategies in large-scale settings, using both theoretical approximations and empirical diagnostics to study gradient flow and the stability of learned representations.
@@
We provide different perspectives on normalization architecture, by casting differently normalized Transformers as variations of a common interacting-particle ODE, where the normalization method determines a speed factor, which can amplify initial velocity and dampen representation collapse in deep layers. Within this unified framework, we extend the framework of mathpersp23 for Post-LN and in particular, establish asymptotic clustering under general conditions on the speed regulation mechanism. To differentiate various normalization schemes we further study the initial and final velocity of tokens corresponding to first and deep layers respectively. In particular, we recover the representation collapse phenomenon that plagues Post-LN. Our theoretical framework identifies Peri-LN as a particularly effective scheme that makes good use of both early and deep layers.
@@ 2.1  Attention
Throughout this paper, we focus on the simplified setting of mathpersp23 where $Q^{t}=K^{t}=V^{t}=I_{d}$ for all $t\geq 0$.
@@
The Root Mean Squared (RMS) norm $\mathsf{Norm}(x)=x/\|x\|$ of a token zhang2019rmsnorm is a critical ingredient of all normalization schemes considered here.
@@
In this paper, we study six major schemes: Post-LN (vaswani2017attention), Pre-LN (xiong2020layer), Mix-LN (li2024mixln), Peri-LN (kim2025peri), nGPT (loshchilov2024ngpt), and sqrt-scaling (sun2025curse). Note that Mix-LN is a combination of Post-LN for $t\leq\tau$ and Pre-LN for $t>\tau$ while sqrt-scalingis a deterministic rescaling of Post-LN. The four remaining schemes are presented in Figure 1. Such explicitly layer-normalization rules are not the only strategies employed in practice. Other attempts to improve normalization suggest better initializations $Q^{0},K^{0},V^{0}$ (kedia2024transformersstableendtoendsignal) and explicit scaling of the updates, similarly to $\alpha_{t}$ that is trainable in nGPT.
@@
sqrt-scaling $X^{t+1}=\mathsf{Norm}\bigl(X^{t}+\frac{1}{\sqrt{t+1}}A^{t}(X^{t})\bigr)$ $\dot{X}(t)=\frac{1}{\sqrt{t+1}}\mathbf{P}_{X(t)}A^{t}(X(t))$
@@ 2.3  Speed regulation formulation
In Post-LN, nGPT, and sqrt-scaling, tokens are constrained to the the sphere ${\mathcal{S}}^{d-1}$ with sqrt-scaling simply adjusting the speed of the particles as a function of $t$ compared to Post-LN. For the other rules where tokens may have varying magnitude, one final projection is typically applied before the final decoding layer in practice. In particular, this means that decoding depends on directions $\theta_{j}(t)=\mathsf{Norm}(x_{j}(t))$, for $j=1,\ldots,n$.
@@ 3  Asymptotic clustering
Since the work of geshkovski2023emergence; mathpersp23, theoretical analyses of attention dynamics have primarily focused on establishing asymptotic clustering under the Post-LN scheme, namely $\theta_{j}(t)\to\theta^{*}$ as $t\to\infty$ for all $j=1,\ldots,n$, under a generic initialization; see also boumal; CheLinPol25. However, empirical studies have revealed that in practice, tokens often remain trapped in metastable states for extended periods before clustering emerges (GesKouPol24; BruPasAga24). Despite this, the clustering phenomenon appears to occur at multiple local scales, and the simplified setting considered in prior work continues to offer valuable insights, as we will demonstrate in the next section. In this section, we extend the analysis and show that asymptotic clustering persists beyond the original Post-LN framework to other normalization schemes.
@@
The main observation of mathpersp23 is that when $KQ^{\top}=QK^{\top}=V$, the Post-LN system is a gradient flow for the energy function
@@
Consider the normalized attention dynamics (NA) with $Q=K=V=I_{d}$. Then for uniformly sampled initializations $\Theta(0)\in({\mathcal{S}}^{d-1})^{\otimes n}$ Post-LN, nGPT, sqrt-scaling cluster asymptotically
@@
In fact, this result holds not only for $Q^{t}=K^{t}=V^{t}=I_{d}$ but more generally for $Q^{t}=Q,K^{t}=K$, and $V^{t}=V=Q^{\top}K=K^{\top}Q$ as in sander2022sinkformers. The second condition on the magnitude growth can be traced with Table 2 definition to work with further. For example, we immediately get the following.
@@ 4.1  Prelude: Symmetric initialization
Following mathpersp23; cowsik2024geometric, we begin with a so-called orthogonal symmetric initialization where $\langle\theta_{j}(0),\theta_{k}(0)\rangle=0$ for $j\neq k$ and $r_{j}(0)=1$ for all $j$. This configuration approximately matches that of randomly initialized tokens in high dimension. Due to the symmetry, the cosine similarity $\gamma(t)=\langle\theta_{j}(t),\theta_{k}(t)\rangle$ does not depend on $j\neq k$ and the entire token dynamics reduces to the evolution of two scalar quantities: $\gamma(t)$ and $r(t)$. In the Appendix, we derive a simple ODE for $\gamma(t),r(t)$ following mathpersp23. We plot ODE-based evolution of $\gamma(t)$ in Figure 2 with parameters $\beta=5$, $n=256$. Despite its simplicity, this setup already provides striking insight into the effects of different normalization schemes. The importance tracks in how cosine similarity evolution is alike in the theoretical formula plotted in Figure 2 and the experimental setup with random weights modeled in Figure 4.
@@
sqrt-scaling $\frac{2}{e^{\beta}+n-1}$ $C\frac{e^{-4\sqrt{t}}}{\sqrt{t}}$
@@
A few remarks are in order. First, the initial velocities are comparable across models, up to the effects of the tuning parameters $\alpha_{0}$ and $r_{0}$. Notably, the temperature parameter $\beta$ exponentially damps the initial velocity, suggesting that initializing $Q$ and $K$ with smaller magnitudes in the early layers may be beneficial. More striking is the effect of speed regulation at terminal velocity: Pre-LN, Mix-LN, and nGPT (with constant $\alpha_{t}$) exhibit a polynomial slowdown, in contrast to other normalization schemes. While sqrt-scaling converges more slowly than exponential, it still outpaces the polynomial decay. This implies that Pre-LN, Mix-LN, and nGPT cluster more gradually than their counterparts—indicating a more effective use of intermediate layers and a stronger resistance to representation collapse. Finally, note that the trainable parameter $\alpha_{t}$ in nGPT can have a drastic impact on both initial and terminal velocity. See Figure 2 for a visual representation of cosine similarity and evolution of $\dot{\gamma}$ relative to time and position. See Figure 3 for comparison between different $\alpha_{t}$ in nGPT.
@@ 4.2  Initial velocity
To interpret the significance of Theorem 4.2, recall from Table 2 that the initial velocity of direction $\theta_{j}$ is dampened by a factor proportional to $\|A_{j}(0)\|$ for both Peri-LN and nGPT. Consequently the first–layer angular displacement of Peri-LN and nGPT exceeds that of Post-LN, Pre-LN, Mix-LN, and sqrt-scaling, by a factor $\Omega(\min(d/\log n,\sqrt{n/\log n}))$.
@@ 4.3  Terminal velocity
Radial Growth under Pre-LN. Our first goal is to estimate the rate of growth of $r_{j}(t)$, the norm of token $j$’s representation, under Pre-LN normalization. Empirically, the growth of hidden states in transformers has been well-documented. For instance, studies such as (xiong2020layer; kedia2024transformersstableendtoendsignal) observe that in randomly initialized transformers, $r_{j}(t)\sim\sqrt{t}$, reflecting the diffusive nature of a random walk induced by randomly sampled projections $V$.
@@
Again this result corroborates the findings of Section 4.1: Post-LN and sqrt-scaling cluster token directions at an exponential rate, while Pre-LN, Peri-LN and Mix-LN slow down to a polynomial ($\sim 1/t^{C}$) decay. Moreover, nGPT has the ability to control rate of clustering through $\alpha_{t}$. This confirms that Pre-LN makes better use of depth, as tokens continue to evolve meaningfully across many layers, rather than collapsing too quickly. In particular, this difference explains why Pre-LN is less prone to representation collapse in very deep models compared to Post-LN.